# A joint optimization approach to identifying sparse dynamics using least squares kernel collocation

### v1

@@ 2.2 Discretization
which we use to compute our estimators. At this point, the reproducing property and boundedness of derivative evaluations in the RKHS $\mathcal{X}$ give a natural parametrization of $\widehat{\boldsymbol{x}}$ which is justified by the following representer theorem, which converts optimization over the infinite-dimensional space $\mathcal{X}$ into a problem on a finite-dimensional subspace. This result is more general than the typical representer theorems often found in ML, such as in [Rasmussen2006Gaussian], which only make use of pointwise evaluations, but is very similar to the representer theorem given in [CHEN2021110668], and follows from the more general results of [Owhadi2019-xh]. The purpose of this result is to reduce the optimization problem over the infinite-dimensional space $\mathcal{X}$ into an optimization problem over the finite-dimensional subspace spanned by an explicit set of basis functions.
@@
Recall that linear operators of the form $\boldsymbol{x}\mapsto\frac{d^{r}}{dt^{r}}\boldsymbol{x}(\tau)$ for r=0,,pr=0,...,p and $\tau\in[0,T]$ are assumed to be bounded on the RKHS $\mathcal{X}$. Thus, by the reproducing property [ZhouDerivative],
@@
which gives the structure of the matrix $\boldsymbol{K}$, and the RKHS norm of individual component of the trajectory $x_{i}(\cdot)\in\operatorname{span}\left\{\partial_{2}^{r}k(\cdot,\tau_{m}),\,\,r=0,1,...,p,m=1,...,M\right\}$. Stacking the individual components together and noting that
@@
Consider the optimization problem (8). Let $\alpha,\beta,\lambda,\mu>0$. Let $\mathcal{X}$ be the vector-valued RKHS associated to the $d$-fold Cartesian product of an RKHS $\mathcal{H}_{k}$ associated to the positive definite kernel $k$. Assume that $\mathcal{H}_{k}$ is compactly embedded in $C^{p}([0,T],\mathbb{R}^{d})$. Assume that all functions $\phi_{q}$ in the library $\Phi$ are continuous on $\mathbb{R}^{dp}$. Then there exists a minimizer $(\boldsymbol{x}^{*},\boldsymbol{\theta}^{*})\in\mathcal{X}\times\mathbb{R}^{d\times Q}$
@@
Because the features $\phi_{q}$ are assumed to be continuous functions, and point evaluations of derivatives up to order $p$ are bounded on $\mathcal{H}_{k}$ due to the compact embedding property, $\mathcal{L}_{C}$ and $\mathcal{L}_{D}$ are continuous on both $\mathcal{X}\times\mathbb{R}^{d\times Q}$ and $C^{p}([0,T]\times\mathbb{R}^{d})$. Additionally, both are lower bounded by zero. Because $R$ is lower semicontinuous on $\mathbb{R}^{d\times Q}$, we have $\mathcal{L}$ is lower semicontinuous on $C^{p}([0,T],\mathbb{R}^{d})$. Since $\frac{\lambda}{2}\|\boldsymbol{x}\|_{\mathcal{X}}^{2}+\frac{\mu}{2}\|\boldsymbol{\theta}\|_{F}^{2}$ is coercive on $\mathcal{X}\times\mathbb{R}^{d\times Q}$, we obtain that $\mathcal{L}$ is coercive as well. For a minimizing sequence $(\boldsymbol{x}^{(k)},\boldsymbol{\theta}^{(k)})$, and we may extract a subsequence that strongly converges in $(C^{p}([0,T],\mathbb{R}^{d})$ to a candidate minimizer $(\boldsymbol{x}^{*},\boldsymbol{\theta}^{*})$. Lower semicontinuity of $\mathcal{L}$ yields that $\lim_{k\to\infty}\mathcal{L}(\boldsymbol{x}^{(k)},\boldsymbol{\theta}^{(k)})=\mathcal{L}(\boldsymbol{x}^{*},\boldsymbol{\theta}^{*})$. ∎
@@ 3 Algorithms and implementation details
In order to numerically compute solutions to 26, we alternate between applying a Levenberg-Marquardt (LM) algorithm to solve smooth nonlinear least squares problems with the support of $\boldsymbol{\theta}$ restricted, and a sparse regression algorithm to choose the support of $\boldsymbol{\theta}$. We first write $\mathcal{J}$ as a regularized nonlinear least squares objective. Define the residual functions
@@ 3.1 Optimization
    $(\boldsymbol{z}^{(k+1)},\boldsymbol{\theta}^{(k+1)})=\operatorname*{arg\,min}_{\boldsymbol{z},\boldsymbol{\theta}}\mathcal{J}(\boldsymbol{x},\boldsymbol{\theta})\text{, subject to }\,\mathrm{supp}(\boldsymbol{\theta})\subset S^{(k)}$ (30)
@@
We initialize with a dense support by setting $S^{(0)}$ to be all available indices of $\boldsymbol{\theta}$, and alternate these two steps until the support $S^{(k)}$ stabilizes, i.e. $S^{(k)}=S^{(k-1)}$, which corresponds to a fixed point. In most of our examples, we take $\mathcal{S}$ to be the result of sequentially thresholded ridge regression [Brunton2016], which is also the choice taken in [ijcai2021p283, Chen2021] and makes our algorithm template match their ADO approach. Options for sparsifiers are discussed further in Section 3.3.
@@ 3.2 Smooth Nonlinear Least Squares
To solve the smooth subproblems arising in the active-set refinement eq. 30, we use an LM method, which is a standard and well studied algorithm for nonlinear least squares[levenberg, marquardt, Aravkin2024-bz, wright1999numerical]. For notational convenience, stack the variables $\boldsymbol{z},\boldsymbol{\theta}$ and restrict to the current support:
@@
If $\rho^{(k)}<0.01$, we reject the step and increase $\gamma^{(k)}$; otherwise we accept and update $\gamma^{(k)}$ for the next iteration. One practical choice, analyzed in [Aravkin2024-bz], is
@@ 3.3 Sparsifying iterations
As mentioned in Section 3.1, we alternate between solving the fixed support, smooth nonlinear least squares problem (30) and updating the support of the coefficient set to obtain $S^{k+1}$, applying a sparse feature selection method in eq. 31. Hence, a crucial component of our algorithm is the choice of regularizer or feature selection procedure in eq. 31. The work that introduced SINDy [Brunton2016] suggested ridge regularized sequentially thresholded least squares (STLSQ). This alternates between solving an ridge regularized least squares problem, and evicting features whose coefficients fall below a given threshold. Another approach is to add regularization which promotes sparsity. These include LASSO [tibshirani1996lasso] which regularizes with an $\ell_{1}$ penalty, SR3[SR3, Champion2020] which minimizes smoothed versions of nonsmooth penalties, and MIOSR [bertsimas_learning_2023], which directly solves $\ell_{0}$ constrained problems using mixed integer programming. Ensembling the approaches above via bagging and feature subsampling can sometimes improve the performance of such algorithms [fasel_ensemble-sindy_2022], and provide some uncertainty quantification [ensemble_uncertainty].
An alternative approach to feature selection is to define an appropriate sparsity promoting prior distribution and likelihood, and approximately sample from the posterior. These include spike-and-slab priors [Gao2022] and the regularized horseshoe prior [bayesiansindy].
In most of our experiments, we apply ridge regularized STLSQ as it empirically performed the best, and was easy to tune. In section 4.5.2 we applied SR3 [SR3] with an L1 penalty–we found this to be more consistent on varying trajectory lengths and noise levels without any retuning, but it can difficult to intuit good parameters a priori. In certain cases, including ensembling within our sparsification procedure greatly improved the robustness of our method, see section 4.5.1. At the cost of losing a straightforward variational problem, considering an algorithmic sparsification framework allows the specific choice of sparsifier to become an implementation detail that is easy for users to modify and provides substantial flexibility.
@@ 4 Numerical Results
We validate JSINDy on a range of examples that illustrate its capability in handling noisy, scarce and partial observations, higher order systems, and model misspecification.
@@
This is not relevant in Section 4.2 due to unidentifiability, and in Section 4.4, as there is no “true” coefficient matrix. All models, for each experiment, use a sum of a scaled constant kernel and Matèrn kernel with $\nu=11/2$, which corresponds to five-times differentiable functions.
We test our algorithm on a variety of systems and observation patterns. We start with the Lorenz 63 system, which is commonly used for testing methods for identifying nonlinear dynamics, using a relatively standard sampling frequency and noise level. Next, we consider the regime of scarce sampling, with sampling frequencies far below what have been previously considered in the literature, working with the Lorenz and Lotke–Volterra dynamics. We then consider problems with partial observations, where we never observe the entire state at once (even with noise), with cases of alternating between observing the different coordinates, and observing only a subset. Finally, we consider the related problem of identifying higher order ODEs from only observations of the state variable, but not its derivatives, working with the Van der Pol system, a higher dimensional coupled version of the Duffing system, and the nonlinear pendulum under model miss-specification. The problem of identifying higher order ODEs from point observations can can be seen as a version of a partially observed problem with additional structure between dimensions if reduced to first order.
In all of these examples, we find that the assumption that $\boldsymbol{x}(t)$ satisfies some autonomous ODE is a strong prior that greatly improves inference, even when the ODE must itself be identified. Identification of the ODE itself is further facilitated by the assumption of sparsity in the dynamics which significantly restricts the class of models under consideration. The benefits of a joint approach to learning as opposed to two-step estimation have been observed in many works [Fung2025-odr, Chen2021, ijcai2021p283, Hokanson2023-sr, Kaheman_2022, jalalian2025dataefficientkernelmethodslearning]. The problem of data scarcity was considered in [Chen2021, jalalian2025dataefficientkernelmethodslearning, Hokanson2023-sr]. The SIDDS algorithm of [Hokanson2023-sr] found that such an approach can reach the Cramer-Rao lower bound and demonstrated the necessity of decoupling the discretization of the ODE from the grid of observations in order to reduce bias when the sampling period is large. Our work expands the flexibility to even scarcer data and partial observations.
@@
Figure 1: Lorenz 63 system with observations sampled at $\Delta t=0.05$ until $t=10$ with added noise set to $\sigma^{2}=4$. Simulated dynamics from LABEL:eqn:lorenz are compared against true dynamics from $t=10$ until $t=15$.
@@ 4.1 Scarce observations
We now consider cases of very scarcely sampled data with the Lorenz 63 and the Lotka-Volterra systems.
@@ 4.1.1 Lorenz 63 Scarce
with $\operatorname{\operatorname{RMSE}_{\text{Var}}(\widehat{\boldsymbol{x}},\boldsymbol{x})}=1.608\cdot 10^{-3}$ and $\operatorname{\operatorname{RMSE}_{\text{Fro}}(\widehat{\boldsymbol{\theta}},\boldsymbol{\theta})}=2.356\cdot 10^{-3}$. These results are very accurate, owing to the lack of noise. Indeed, a parameter counting argument suggests that once the support is identified, this problem is essnetially a noiseless and well-posed though ill-conditioned inversion problem, where we should expect good algorithms to attain very high accuracy.
@@
4.1.2 Lotka Volterra
The Lotka-Volterra system is a nonlinear oscillator with periodic solutions, describing population dynamics of a predator-prey ecology, and is a common benchmark for system identification[bertsimas_learning_2023, fasel_ensemble-sindy_2022, Kaheman_2022, haas2024kalman]. This system governs the populations of prey $x_{1}$ and predators $y_{1}$:
$\begin{split}\frac{dx_{1}}{dt}&=p_{11}x_{1}-p_{12}x_{1}x_{2},\\ \frac{dx_{2}}{dt}&=-p_{21}x_{2}+p_{22}x_{1}x_{2},\end{split}$ (43)
with parameters $p_{11}=p_{12}=p_{21}=1.$, and $p_{22}=0.5$. We again considered scarce sampling, with approximately $2$ observations per cycle, taken at a sampling period of $\Delta t=3$ in the interval $t\in[0,50]$. Gaussian noise with variance $\sigma^{2}=0.04$ was added to the observations. We set the feature library to include monomials up to second order:
@@ 4.1.2 Lotka Volterra
Figure 3: Results for the Lotke-Volterra system of section 4.1.2. Samples are taken at the rate $\Delta t=3$ up to $t=48$, including additive noise with $\sigma=0.2$. Simulated dynamics are shown up to $t=100$, comparing the using the results from (44) with the true dynamics.
Figure 4: Left: True Lotka-Volterra plotted from multiple initial conditions on phase portrait using (43). Red trajectory and measurements are those from Figure 3. Right: Using learned model (44) to simulate same trajectories provided same initial conditions.
@@ 4.2.1 Alternating coordinates
In the first case, we alternated between periods of observing each of the three coordinates one at a time for $10$ time steps each. For each coordinate, this leads to relatively long gaps (of duration $t=0.5$ ) where no observations of the state are made. However, our approach allows the dynamics to effectively fill in this missing information, correlating the data available between the different coordinates. Notably, if the dynamics were known ahead of time, then applying a high order nonlinear Kalman smoothing method could be similarly effective, but this is not applicable here, since there is no a priori knowledge of the governing equations [Stevens-Haas2021-mn]. Our model learned the system in (45) with $\operatorname{\operatorname{RMSE}_{\text{Fro}}(\widehat{\boldsymbol{\theta}},\boldsymbol{\theta})}=5.692\cdot 10^{-3}$ and $\operatorname{\operatorname{RMSE}_{\text{Var}}(\widehat{\boldsymbol{x}},\boldsymbol{x})}=1.961\cdot 10^{-2}$. Results can be seen in Figure 5.
@@
Figure 5: Results for Lorenz 63 system with streams of partial observations. Observations alternative between each coordinate every 10 samples at sampling rate $\Delta t=0.025$ up to $t=10$ with added noise set to $\sigma^{2}=0.1$. Learned dynamics are simulated from $t=10$ until $t=20$.

### v2

@@ 2.2 Discretization
which we use to compute our estimators. At this point, the reproducing property and boundedness of derivative evaluations in the RKHS $\mathcal{X}$ give a natural parametrization of $\widehat{\boldsymbol{x}}$ which is justified by the following representer theorem, which converts optimization over the infinite-dimensional space $\mathcal{X}$ into a problem on a finite-dimensional subspace. This result is more general than the typical representer theorems often found in ML, such as in [74], which only make use of pointwise evaluations, but is very similar to the representer theorem given in [21], and follows from the more general results of [68]. The purpose of this result is to reduce the optimization problem over the infinite-dimensional space $\mathcal{X}$ into an optimization problem over the finite-dimensional subspace spanned by an explicit set of basis functions.
@@
Recall that linear operators of the form $\boldsymbol{x}\mapsto\frac{d^{r}}{dt^{r}}\boldsymbol{x}(\tau)$ for r=0,,pr=0,...,p and $\tau\in[0,T]$ are assumed to be bounded on the RKHS $\mathcal{X}$. Thus, by the reproducing property [104],
@@
which gives the structure of the matrix $\boldsymbol{K}$, and the RKHS norm of an individual component of the trajectory $x_{i}(\cdot)\in\operatorname{span}\left\{\partial_{2}^{r}k(\cdot,\tau_{m}),\,\,r=0,1,...,p,m=1,...,M\right\}$. Stacking the individual components together and noting that
@@
Consider the optimization problem (8). Let $\alpha,\beta,\lambda,\mu>0$. Let $\mathcal{X}$ be the vector-valued RKHS associated to the $d$-fold Cartesian product of an RKHS $\mathcal{H}_{k}$ associated to the positive definite kernel $k$. Assume that $\mathcal{H}_{k}$ is compactly embedded in $C^{p}([0,T],\mathbb{R}^{d})$. Assume that all functions $\phi_{q}$ in the library $\Phi$ are continuous on $\mathbb{R}^{dp}$. Then there exists a minimizer $(\boldsymbol{x}^{*},\boldsymbol{\theta}^{*})\in\mathcal{X}\times\mathbb{R}^{d\times Q}$.
@@
Because the features $\phi_{q}$ are assumed to be continuous functions, and point evaluations of derivatives up to order $p$ are bounded on $\mathcal{H}_{k}$ due to the compact embedding property, $\mathcal{L}_{C}$ and $\mathcal{L}_{D}$ are continuous on both $\mathcal{X}\times\mathbb{R}^{d\times Q}$ and $C^{p}([0,T],\mathbb{R}^{d})\times\mathbb{R}^{d\times Q}$. Additionally, both are lower bounded by zero. Since $\frac{\lambda}{2}\|\boldsymbol{x}\|_{\mathcal{X}}^{2}+\frac{\mu}{2}\|\boldsymbol{\theta}\|_{F}^{2}$ is coercive on $\mathcal{X}\times\mathbb{R}^{d\times Q}$, we obtain that $\mathcal{L}$ is coercive as well. For a minimizing sequence $(\boldsymbol{x}^{(k)},\boldsymbol{\theta}^{(k)})$, we may extract a subsequence that strongly converges in $C^{p}([0,T],\mathbb{R}^{d})\times\mathbb{R}^{d\times Q}$ to a candidate minimizer $(\boldsymbol{x}^{*},\boldsymbol{\theta}^{*})$. Lower semicontinuity of $\mathcal{L}$ yields that $\lim_{k\to\infty}\mathcal{L}(\boldsymbol{x}^{(k)},\boldsymbol{\theta}^{(k)})=\mathcal{L}(\boldsymbol{x}^{*},\boldsymbol{\theta}^{*})$ giving existence of a minimizer. ∎
@@ 3 Algorithms and implementation details
In order to numerically compute solutions to eq. 26, we alternate between applying a Levenberg-Marquardt (LM) algorithm to solve smooth nonlinear least squares problems with the support of $\boldsymbol{\theta}$ restricted, and a sparse regression algorithm to choose the support of $\boldsymbol{\theta}$. We first write $\mathcal{J}$ as a regularized nonlinear least squares objective. Define the residual functions
@@ 3.1 Optimization
    $(\boldsymbol{z}^{(k+1)},\boldsymbol{\theta}^{(k+1)})=\operatorname*{arg\,min}_{\boldsymbol{z},\boldsymbol{\theta}}\mathcal{J}(\boldsymbol{z},\boldsymbol{\theta})\text{, subject to }\,\mathrm{supp}(\boldsymbol{\theta})\subset S^{(k)}$ (30)
@@
We initialize with a dense support by setting $S^{(0)}$ to be all available indices of $\boldsymbol{\theta}$, and alternate these two steps until the support $S^{(k)}$ stabilizes, i.e. $S^{(k)}=S^{(k-1)}$, which corresponds to a fixed point. In most of our examples, we take $\mathcal{S}$ to be the result of sequentially thresholded ridge regression [12], which is also the choice taken in [87, 22] and makes our algorithm template match their ADO approach. Options for sparsifiers are discussed further in Section 3.3.
@@ 3.2 Smooth nonlinear least squares
To solve the smooth subproblems arising in the active-set refinement eq. 30, we use an LM method, which is a standard and well studied algorithm for nonlinear least squares [54, 59, 2, 99]. For notational convenience, stack the variables $\boldsymbol{z},\boldsymbol{\theta}$ and restrict to the current support:
@@
If $\rho^{(k)}<0.01$, we reject the step and increase $\gamma^{(k)}$; otherwise we accept and update $\gamma^{(k)}$ for the next iteration. One practical choice, analyzed in [2], is
@@ 3.3 Sparsifying iterations
As mentioned in Section 3.1, we alternate between solving the fixed support, smooth nonlinear least squares problem (30) and updating the support of the coefficient set to obtain $S^{k+1}$, applying a sparse feature selection method in eq. 31. Hence, a crucial component of our algorithm is the choice of regularizer or feature selection procedure in eq. 31. The work that introduced SINDy [12] suggested ridge regularized sequentially thresholded least squares (STLSQ). This alternates between solving a ridge regularized least squares problem, and evicting features whose coefficients fall below a given threshold. Another approach is to add regularization which promotes sparsity. These include LASSO [91] which regularizes with an $\ell_{1}$ penalty, SR3[103, 18] which minimizes smoothed versions of nonsmooth penalties, and MIOSR [8], which directly solves $\ell_{0}$ constrained problems using mixed integer programming. Ensembling the approaches above via bagging and feature subsampling can sometimes improve the performance of such algorithms [26], and provide some uncertainty quantification [29].
An alternative approach to feature selection is to define an appropriate sparsity promoting prior distribution and likelihood, and approximately sample from the posterior. These include spike-and-slab priors [60] and the regularized horseshoe prior [37].
In most of our experiments, we apply ridge regularized STLSQ as it empirically performed the best, and was easy to tune. In section 4.5.2 we applied SR3 [103] with an L1 penalty–we found this to be more consistent on varying trajectory lengths and noise levels without any retuning, but it can be difficult to intuit good parameters a priori. In certain cases, including ensembling within our sparsification procedure greatly improved the robustness of our method, see section 4.5.1. At the cost of losing a straightforward variational problem, considering an algorithmic sparsification framework allows the specific choice of sparsifier to become an implementation detail that is easy for users to modify and provides substantial flexibility.
@@ 4 Numerical results
We validate JSINDy on a range of examples that illustrate its capability in handling noisy, scarce and partial observations, higher order systems, and model misspecification. [^1] An implementation of JSINDy and the code used to generate the experiments in this section are available at https://github.com/AHsu98/jsindy.
@@
This is not relevant in Section 4.2 due to unidentifiability, and in Section 4.4, as there is no “true” coefficient matrix. All models, for each experiment, use a sum of a scaled constant kernel and Matérn kernel with $\nu=11/2$, which corresponds to five times differentiable functions.
We test our algorithm on a variety of systems and observation patterns. We start with the Lorenz 63 system, which is commonly used for testing methods for identifying nonlinear dynamics, using a relatively standard sampling frequency and noise level. Next, we consider the regime of scarce sampling, with sampling frequencies far below what have been previously considered in the literature, working with the Lorenz and Lotka–Volterra dynamics. We then consider problems with partial observations, where we never observe the entire state at once (even with noise), with cases of alternating between observing the different coordinates, and observing only a subset. Finally, we consider the related problem of identifying higher order ODEs from only observations of the state variable, but not its derivatives, working with the Van der Pol system, a higher dimensional coupled version of the Duffing system, and the nonlinear pendulum under model misspecification. The problem of identifying higher order ODEs from point observations can be seen as a version of a partially observed problem with additional structure between dimensions if reduced to first order.
In all of these examples, we find that the assumption that $\boldsymbol{x}(t)$ satisfies some autonomous ODE is a strong prior that greatly improves inference, even when the ODE must itself be identified. Identification of the ODE itself is further facilitated by the assumption of sparsity in the dynamics which significantly restricts the class of models under consideration. The benefits of a joint approach to learning as opposed to two-step estimation have been observed in many works [28, 22, 87, 38, 43, 40]. The problem of data scarcity was considered in [22, 40, 38]. The SIDDS algorithm of [38] found that such an approach can reach the Cramer-Rao lower bound and demonstrated the necessity of decoupling the discretization of the ODE from the grid of observations in order to reduce bias when the sampling period is large. Our work expands the flexibility to even scarcer data and partial observations.
@@
Figure 1: Results for Lorenz 63 system. Observations sampled at $\Delta t=0.05$ until $t=10$ with added Gaussian noise of variance $\sigma^{2}=4$. Simulated dynamics from LABEL:eqn:lorenz are compared against true dynamics from $t=10$ until $t=15$.
@@ 4.1 Scarce observations
We now consider cases of very scarcely sampled data with the Lorenz 63 and the Lotka–Volterra systems.
@@ 4.1.1 Lorenz 63 Scarce
with $\operatorname{\operatorname{RMSE}_{\text{Var}}(\widehat{\boldsymbol{x}},\boldsymbol{x})}=1.608\cdot 10^{-3}$ and $\operatorname{\operatorname{RMSE}_{\text{Fro}}(\widehat{\boldsymbol{\theta}},\boldsymbol{\theta})}=2.356\cdot 10^{-3}$. These results are very accurate, owing to the lack of noise. Indeed, a parameter counting argument suggests that once the support is identified, this problem is essentially a noiseless and well-posed though ill-conditioned inversion problem, where we should expect good algorithms to attain very high accuracy.
@@
4.1.2 Lotka–Volterra
The Lotka–Volterra system is a nonlinear oscillator with periodic solutions, describing population dynamics of a predator-prey ecology, and is a common benchmark for system identification [8, 26, 43, 86]. This system governs the populations of prey $x_{1}$ and predators $x_{2}$:
$\begin{split}\frac{dx_{1}}{dt}&=x_{1}-x_{1}x_{2},\\ \frac{dx_{2}}{dt}&=-x_{2}+x_{1}x_{2},\end{split}$ (43)
We again considered scarce sampling, with approximately $2$ observations per cycle, taken at a sampling period of $\Delta t=3$ in the interval $t\in[0,50]$. Gaussian noise with variance $\sigma^{2}=0.04$ was added to the observations. We set the feature library to include monomials up to second order:
@@ 4.1.2 Lotka–Volterra
Figure 3: Results for the Lotka–Volterra system of section 4.1.2. Samples are taken at the rate $\Delta t=3$ up to $t=48$, including additive noise with $\sigma=0.2$. Simulated dynamics are shown up to $t=100$, comparing the results from (44) with the true dynamics.
Figure 4: Left: True Lotka–Volterra plotted from multiple initial conditions on phase portrait using (43). Red trajectory and measurements are those from Figure 3. Right: Using learned model (44) to simulate same trajectories provided same initial conditions.
@@ 4.2.1 Alternating coordinates
In the first case, we alternated between periods of observing each of the three coordinates one at a time for $10$ time steps each. For each coordinate, this leads to relatively long gaps (of duration $t=0.5$ ) where no observations of the state are made. However, our approach allows the dynamics to effectively fill in this missing information, correlating the data available between the different coordinates. Notably, if the dynamics were known ahead of time, then applying a high order nonlinear Kalman smoothing method could be similarly effective, but this is not applicable here, since there is no a priori knowledge of the governing equations [7, 80]. Our model learned the system in (45) with $\operatorname{\operatorname{RMSE}_{\text{Fro}}(\widehat{\boldsymbol{\theta}},\boldsymbol{\theta})}=5.692\cdot 10^{-3}$ and $\operatorname{\operatorname{RMSE}_{\text{Var}}(\widehat{\boldsymbol{x}},\boldsymbol{x})}=1.961\cdot 10^{-2}$. Results can be seen in Figure 5.
@@
Figure 5: Results for Lorenz 63 system with streams of partial observations. Observations alternate between each coordinate every 10 samples at sampling rate $\Delta t=0.025$ up to $t=10$ with added noise set to $\sigma^{2}=0.1$. Learned dynamics are simulated from $t=10$ until $t=20$.